Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2015-10-10

Observers: Joe Lyman, Heather Campbell
TAT: Janet Chen, Annalisa de Cia
Data reducers: Morgan Fraser

Intermittent thick clouds for first half of night, causing problems for a few exposures.
Conditions improved for second half.

UT      Target                  Comment
========================================================================
23:30                           Cloudy, wind ~5m/s, dome closed.
00:45                           Dome opened, intermittent clouds
01:00   MASTERJ233...           Classification (guide star lost and aborted after ~400s, noisy but looks to be type II)
01:20   PS15cgd                 Classification (aborted, nothing visible in 2D spectrum)
01:35                           Dome closed.
02:45                           Dome opened
02:45   PS15cgd                 Classification reattempt (blue continuum with gal/agn features)
03:20   ASASSN-15lh             Followup (g11)
03:20                           Seeing ~1.2-1.3"
04:25   SN2009ip                Followup photometry (BVRi)
06:00   DES15S2nr               Followup (g13)
07:50   ASASSN-15og             Followup photometry (BVRi)
08:00   SN2015F                 Followup photometry (UBVgri)
08:20                           Seeing ~1.5"
08:35   L745a                   Spec standard
09:00   PS15cej                 Classification (Ia at peak)
09:25                           Sky flats UVi

Comments